OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!!!!!!!!!!!! Which expression could be used to solve the problem? While preparing a recipe, Chris noticed that 4 cans held a total of 34 oz of tomato sauce. The cans were all the same size. How many ounces did each can hold?

OpenStudy (anonymous):

A. 34*4 B.34+4 C.\[34\div4\] D.\[4\div34\]

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@Catlover5925

OpenStudy (anonymous):

@CarlosJose

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

@chosenmatt

OpenStudy (chosenmatt):

what do you think?

OpenStudy (chosenmatt):

@CausticSyndicalist

OpenStudy (anonymous):

help

OpenStudy (anonymous):

@chosenmatt I don't know the answer or could just some one help me its been 21 minutes to know what the answer will be!???!????!?!?/??!?!?!?!!?!?

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

if 4 cans hold 34 oz together, how could we find out how much 1 holds?

OpenStudy (anonymous):

umm by divinding?????

OpenStudy (aripotta):

exactly. so 34/4 is your answer

OpenStudy (anonymous):

can you help me in 3 more please

OpenStudy (aripotta):

i can check your answers for you

OpenStudy (anonymous):

The Quinn family drove 228 miles in 4 hours at a constant rate. Which equation can be used to determine how far they traveled each hour? @AriPotta

OpenStudy (aripotta):

what do you think? it's exactly like the other one

OpenStudy (anonymous):

A. d=228 divided by 4 B. d=228 times 4 C. d=228+4 D. d=228-4

OpenStudy (anonymous):

maybe its B

OpenStudy (anonymous):

or C @AriPotta hello???

OpenStudy (aripotta):

no, it's a

OpenStudy (aripotta):

you have to divide to get the amount for 1

OpenStudy (anonymous):

oh ok can you help with 2 more